OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1 (1, 5) (-5, -1) (5, 1) (-5, 1)

OpenStudy (ivettef365):

you can arrange the x -4y = 1 as x = 4y +1 then substitute the x for 4y +1 on the first equation

OpenStudy (anonymous):

is it A ?

OpenStudy (ivettef365):

lets try it A = (1,5) 2(1) - 9 (5) = 1 2 - 45 = 1 -43 = 1 not true can not be A

OpenStudy (anonymous):

Its C. :)